Two identical firms compete as a Cournot duopoly. The inverse market demand they face is P = 120-2Q. The total cost function for each firm is TC1(Q) = 4Q1. The total cost function for firm 2 is TC2(Q) = 2Q2. What is the output of each firm?

Find:

Q1 = ?

Q2 = ?
Answer #1

Here; p = 120-20 Where Q 9, +2 P = 120-29, -292 for Firmy: TIF 49, MC, a dCT(1) aq, 4 and) TRI= Pixa, = 1209, - 29,2- 202 MRMIR2= d [TR2) dq2 = 120-29, -3492 Now MC2 = MIR2 120 - 24-492 = 2 29, + & 2 = 118 -- tin Now Solving equation ci, and is for